Lemon drops worth $3.80 per pound are to be mixed with jelly beans that cost $2.40 per pound to make 100 pounds of a mixture worth $2.96 per pound. How many pounds of each candy should be used?:
Answer by stanbon(75887)

Lemon drops worth $3.80 per pound are to be mixed with jelly beans that cost $2.40 per pound to make 100 pounds of a mixture worth $2.96 per pound. How many pounds of each candy should be used?:
------------------------
Value equation:
value + value = value
3.80x + 2.40(100-x) = 2.96*100
Multiply thru by 100 to get:
380x + 24000 - 240x = 29600
140x = 1560
x = 40 lbs (amt. 0f $3.80 per pound lemon drops in the mixture)
100-x = 60 lbs (amt. of $240 per pound jelly beans needed in the mixture)
============================================================================
Cheers,
Stan H.